OpenStudy (anonymous):

Over which interval does the function decrease the fastest? when f(x)=3rd power sqreroot(-x+5)

OpenStudy (anonymous):

pleaseeeeeee help anyoneee

OpenStudy (anonymous):

its 330 am its my last math problem I can't get it wrong and I can't solve on my own

OpenStudy (perl):

did you take the derivative twice

OpenStudy (anonymous):

what do you mean, no I don't think so.

OpenStudy (perl):

the second derivative will tell you on which interval it is decreasing the fastest

OpenStudy (perl):

the first derivative will tell you on which interval it is decreasing

OpenStudy (jhannybean):

second derivative also tells you instantaneous velocity at a point. :-)

OpenStudy (anonymous):

Oh ok well I wasn't taught what the second one was hoe do I get it? like how to solve for it?

OpenStudy (jhannybean):

I'm guessing the 4th derivative (imaginary) will tell you when it's decreasing at light speed! aha.

OpenStudy (jhannybean):

Were you taught the first derivative?

OpenStudy (anonymous):

I don't know how to solve this eqaution at all could you tell me what steps to take first?

OpenStudy (jhannybean):

okok, are you taking calculus right now? Or is this just algebra?

OpenStudy (anonymous):

its algebra statistics

OpenStudy (jhannybean):

Ohh.... hmm Im not too sure. @perl

OpenStudy (anonymous):

I think she left thanks for looking @Jhannybean

OpenStudy (jhannybean):

hmm...maybe maybe not!

OpenStudy (perl):

the function is f(x) = (-x+5)^(3/2) , correct?

OpenStudy (perl):

My reasoning was: sqrt(-x + 5)^3 = [(-x+5)^(1/2) ] ^3 = (-x+5)^(3/2 )

OpenStudy (perl):

Note that the function is only defined for x <= 5 , since the expression under the square root must be greater than or equal to zero. f ' (x) = 3/2 * (-x + 5) ^(1/2) * (-1) f ' (x) = -3/2 (-x+5)^(1/2) critical points are x = 5 so f is decreasing for x <= 5 f ' ' (x) = -3/2 (-x + 5) ^(-1/2) (-1) f ' ' (x) = 3/2 (-x+5)^(-1/2) it is concave down for x <= 5 Since f is decreasing and concave down it is decreasing the fastest when x = -infinity There might be a typo in your question, since the answer is kind of strange. Or I misunderstood the question.

OpenStudy (perl):

@sandysparklesblahhhhhhh maybe you could upload the original question , take a screen shot

OpenStudy (anonymous):

second derivative is speed? Where you get that Jhanny?

OpenStudy (anonymous):

The first derivative will tell you the rate it is decreasing. The second derivative will tell you the critical points of the first derivative, so that you can minimize the first derivative.